OpenStudy (anonymous):

calculate dy/dt using the given information. y square root x+1 =12; dx/dt = 8, x=15,y=3

OpenStudy (anonymous):

ok i misunderstood the question it says \(y\sqrt{x+1}=12\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then \(y=\frac{12}{\sqrt{x+1}}\) and \[\frac{dy}{dx}=-\frac{6}{\sqrt{(x+1)^3}}\] then \[\frac{dy}{dt}=\frac{dy}{dx}\times \frac{dx}{dt}\] and you are told \(\frac{dx}{dt}=8\)

OpenStudy (anonymous):

how did you get the 6

OpenStudy (anonymous):

was it from the y=3

OpenStudy (anonymous):

\[\frac{ 12 }{ \sqrt{x+1} } = 12\ \left( x+1 \right) ^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

\[\frac{d \left( 12\left( x+1 \right)^\frac{ 1 }{ 2 } \right)}{dx} = \frac{12}{\frac{1}{2}+1}\times \left( x+1 \right)^\left(\frac{1}{2}+1\right)\]

OpenStudy (anonymous):

what happen to x = 15

OpenStudy (anonymous):

the answer is coming out to be -3/4 and i am still not seeing how that is